Dorman Industries has a new project available that requires an initial investment of $5.6 million. The project will provide unlevered cash flows of $785,000 per year for the next 20 years. The company will finance the project with a debt-to-value ratio of .4. The company’s bonds have a YTM of 6.6 percent. The companies with operations comparable to this project have unlevered betas of 1.26, 1.19, 1.41, and 1.36. The risk-free rate is 3.6 percent, and the market risk premium is 6.8 percent. The company has a tax rate of 34 percent.

What is the NPV of this project? (Enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

please give correct answer
Answer #1

Answer:

Calculation of WACC:

Cost of debt = 6.6%

Average unlevered beta = (1.26 + 1.19 + 1.41 + 1.36) / 4 = 1.305

Levered beta = Unlevered beta * (1 + ((1 - Tax rate) * (Debt / Equity)))

= 1.305 * (1+((1- 34%) * (0.4 / 0.6)))

= 1.8792

Cost of equity = Risk free rate + Beta * Market risk premium

= 3.6% + 1.8792 * 6.8%

= 16.37856%

WACC = Cost of equity * Equity weight + Before tax cost of debt * (1 - Tax rate) * Debt weight

= 16.37856% * 60% + 6.6% * (1- 34%) * 40%

= 11.57%

WACC = 11.57%

NPV:

Annual cash flow for 20 years = $785,000

Initial cash flow = $5,600,000

PV factor of $1 annuity at 11.57% rate of discount = (1 - 1/ (1 + 11.57%) 20) / 11.57%

NPV = Annual cash flow * PV factor of annuity - Initial investment

= 785000 * (1 - 1/ (1 + 11.57%) 20) / 11.57% - 5600000

= $425,183.56

NPV of Project = $425,183.56